# CC-Time: Cross-Model and Cross-Modality Time Series Forecasting

### v1

@@ 4.3 Model Analysis
We replace GPT-2 with several more advanced pre-trained language models, including Flan-T5-250M (Chung et al. 2024), LLaMA-7B (Touvron et al. 2023), and LLaMA-13B. As shown in Table 3, different PLM architectures exhibit varying impacts on prediction performance. Overall, adopting stronger PLMs leads to improved forecasting accuracy in CC-Time, demonstrating its ability to leverage the knowledge of PLMs. Interestingly, we observe that the performance improvement from LLaMA-7B to LLaMA-13B is relatively marginal. This suggests that CC-Time may not heavily rely on the additional capabilities of larger models (e.g., reasoning), but instead primarily benefits from their semantic understanding and representation learning. Further experiments about the effectiveness of PLMs are in Appendix  F.1.
@@
To evaluate the effectiveness of correlation modeling, we select two baselines: iTransformer and GPT4TS, to compare with CC-Time. Inspired by the experiment demonstrations of the channel correlations in iTransformer, we calculate the Pearson Correlation Coefficients, a commonly used metric for approximately assessing correlations, between the predicted series and the future series for each model and normalize these results to analyze correlation modeling. We provide a case visualization in Figure 4. Compared to GPT4TS, a channel-independent method, iTransformer and CC-Time effectively model correlations. Compared to iTransformer, the channel correlations predicted by CC-Time are closer to the correlations from the future series, indicating that the correlation modeling of CC-Time is more effective and comprehensive. These observations also further illustrate the potential of PLMs for channel correlation modeling.
@@
Cross-model Analysis
We analyze the modeling approaches of CC-Time, PLM-based methods, and time-series-specific methods. We use a standard metric: the centered kernel alignment (CKA) (Kornblith et al. 2019) to assess the similarity between features and original data. A higher CKA indicates greater similarity, suggesting that the model learns simpler features, while a lower CKA suggests the model learns more complex features. As shown in Figure 5, time-series-specific methods exhibit high CKA values, while PLM-based methods exhibit low CKA values. This suggests that the features learned by the two types of methods differ significantly. The CKA similarity value of CC-Time is intermediate between the two categories of methods, and it achieves the lowest MSE, indicating that the model effectively captures the appropriate complex features through a cross-model modeling approach. Furthermore, CC-Time achieves the best prediction accuracy on two different types of datasets, demonstrating that this cross-model modeling approach can better adapt to diverse time series.
@@ 5 Conclusion
We propose CC-Time, a Cross-Model and Cross-Modality framework for time series forecasting. CC-Time leverages pre-trained language models to jointly model temporal dependency and channel correlations from both time series and their channel text descriptions. Furthermore, CC-Time proposes the cross-model fusion block to adaptively integrate knowledge from the PLMs and time series models to form a comprehensive modeling of time series. Experiment shows that CC-Time achieves state-of-the-art prediction accuracy in both full-data training and few-shot learning.
@@ Appendix A Channel Descriptions
As shown in Figure 6, for generating semantic descriptions, we construct prompts for each channel: This is <dataset name>from <domain>, including <channel name1>, <channel name2>, …, please describe these channels and their correlations. We just need to input the dataset name, domain name, and the specific channel names into the predefined prompt. Then, by entering it into a large language model like ChatGPT, the semantic descriptions of channels are generated. For example, in the Weather dataset (Wu et al. 2021), the description of the temperature channel (denoted as ”T (degC)” in the dataset) is generated from LLMs as follows: Temperature is a key parameter that describes climate conditions. It is closely related to humidity. As the temperature rises, the air can hold more moisture, affecting humidity. CC-Time can leverage this semantic information to understand the correlation between temperature and humidity in the real world, which helps in better modeling patterns.
@@ Appendix B Datasets
ETT (Electric Transformer Temperature) (Zhou et al. 2021) collected from two different electric transformers, spans from July 2016 to July 2018, and includes 7 channels. ETT is divided into four subsets: ETTh1 and ETTh2, recorded hourly, and ETTm1 and ETTm2, recorded every 15 minutes. (2) Weather (Wu et al. 2021) includes 21 different meteorological indicators that provide comprehensive data on weather conditions. These indicators, such as temperature, barometric pressure, humidity and others, offer a broad perspective on the atmospheric environment. (3) Electricity (Trindade 2015) contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. (4) Traffic (Wu et al. 2021) contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. (5) ZafNoo (Poyatos et al. 2020) is from the Sapflux data project including sap flow measurements and environmental variables. (6) CzeLan (Poyatos et al. 2020) is from the Sapflux data project including sap flow measurements and environmental variables. We split each evaluation dataset into train-validation-test sets and detailed descriptions of evaluation datasets are shown in Table 4. According to the channel description generation method proposed in section 3.2, the corresponding text for datasets is generated without requiring any additional human effort.
@@ Appendix C Baselines
  • S2IP-LLM (Pan et al. 2024) aligns the semantic space of pre-trained PLMs with time series embedding space and enhances the representation of time series using semantic space informed prompting to improve forecasting performance.
  • FSCA (Hu et al. 2025) builds the consistent context through structural alignment, logical alignment, and dual-scale GNNs, enabling PLMs to better understand time series.
  • Time-LLM (Jin et al. 2024a) aligns the time series features into the language feature space through reprogramming techniques, then concatenates them with the text prompt before inputting them into a pre-trained large language model for feature extraction.
  • UniTime (Liu et al. 2024b) designs domain instructions as prompts and uses the language model architecture for multi-source time series pre-training to extract broad knowledge.
  • GPT4TS (Zhou et al. 2023) uses the pre-trained GPT2 as the backbone and adapts it to the time series space for feature extraction by fine-tuning the positional encoding and normalization layers.
  • iTransformer (Liu et al. 2024c) treats the entire set of variables as tokens and uses the Transformer architecture to model the correlations between the entire set of channels.
  • Crossformer (Zhang and Yan 2023) proposes a two-stage attention mechanism (cross-time attention and cross-dimension attention) to model the dynamics of time and the correlations between channels.
  • PatchTST (Nie et al. 2023) employs a patching strategy to preserve local information and uses the Transformer architecture to capture the correlation between different patches for temporal modeling. It also applies a channel-independent strategy on the channel dimension.
  • FEDformer (Zhou et al. 2022) proposes a frequency-enhanced decomposed Transformer architecture to model temporal dynamics from the perspective of frequency.
  • TimesNet (Wu et al. 2023) transforms one-dimensional time series into a two-dimensional structure using Fourier transforms, enabling the modeling of 2D temporal variations to capture multi-periodicity time series patterns.
@@ Appendix D Compared with Time Series Foundation Models
To further validate the prediction performance of CC-Time, we conduct comparative experiments with several state-of-the-art time series foundation models, including Chronos (Ansari et al. 2024), MORIAI (Woo et al. 2024), Timer (Liu et al. 2024e), and TTMs (Ekambaram et al. 2024). The evaluation is performed on ETT and Weather datasets, with input lengths of $\{96,336,512\}$ and output lengths spanning $\{96,192,336,720\}$. The final results are selected based on the optimal performance across the three input lengths. As shown in Table 5, CC-Time achieves state-of-the-art prediction performance in most forecasting scenarios, outperforming the four baseline models. This demonstrates that CC-Time not only exhibits strong generalization capabilities but also effectively leverages knowledge from pre-trained language models, leading to significant improvements in time series forecasting accuracy.
@@ F.1 Effectiveness of Pre-trained Language Models
In the second experiment, refer to (Tan et al. 2024), we further validate the effectiveness of PLMs in CC-Time. Specifically, we focus on three aspects: removing the entire PLM layers of CC-Time (W/O PLM), replacing the PLM layers with a single layer of untrained Attention (LLM2Attn), and replacing the PLM layers with a single layer of untrained Transformer (LLM2Trsf). As shown in Table 7, CC-Time effectively utilizes LLM knowledge to model time series. When comparing W/O PLM with CC-Time, the results indicate that relying solely on basic time series modules like Patching and ReVIN is insufficient, and that effectively exploring the potential of PLM knowledge for time series forecasting is crucial. When comparing LLM2Attn and LLM2Trf with CC-Time, the results indicate that relying on single-layer attention or transformer for feature extraction is inadequate, and that fully leveraging the strengths of PLMs and time-series-specific models to capture time series features comprehensively is essential.

### v3

@@ 4.4 Model Analysis
We replace GPT-2 with several more advanced pre-trained language models, including Flan-T5-250M Chung et al. (2024), LLaMA-7B Touvron et al. (2023), and LLaMA-13B. As shown in Table 3, different PLM architectures exhibit varying impacts on prediction performance. Overall, adopting stronger PLMs leads to improved forecasting accuracy in CC-Time, demonstrating its ability to leverage the knowledge of PLMs. Interestingly, we observe that the performance improvement from LLaMA-7B to LLaMA-13B is relatively marginal. This suggests that CC-Time may not heavily rely on the additional capabilities of larger models (e.g., reasoning), but instead primarily benefits from their semantic understanding and representation learning. Further experiments about the effectiveness of PLMs are in Appendix  F.1.
@@
To evaluate the effectiveness of correlation modeling, we select two baselines: iTransformer and GPT4TS, to compare with CC-Time. Inspired by the experiment demonstrations of the channel correlations in iTransformer, we calculate the Pearson Correlation Coefficients, a commonly used metric for approximately assessing correlations, between the predicted series and the future series for each model and normalize these results to analyze correlation modeling. We provide a case visualization in Figure 4. Compared to GPT4TS, a channel-independent method, iTransformer and CC-Time effectively model correlations. Compared to iTransformer, the channel correlations predicted by CC-Time are closer to the correlations from the future series, indicating that the correlation modeling of CC-Time is more effective and comprehensive. These observations also further illustrate the potential of PLMs for correlation modeling.
@@
Cross-model Analysis
We analyze the modeling approaches of CC-Time, PLM-based methods, and time-series-specific methods. We use a standard metric: the centered kernel alignment (CKA) Kornblith et al. (2019) to assess the similarity between features and original data. A higher CKA indicates greater similarity, suggesting that the model learns simpler features, while a lower CKA suggests the model learns more complex features. As shown in Figure 5, time-series-specific methods exhibit high CKA values, while PLM-based methods exhibit low CKA values. This suggests that the features learned by the two types of methods differ significantly.
@@
The CKA similarity value of CC-Time is intermediate between the two categories of methods, and it achieves the lowest MSE, indicating that the model effectively captures the appropriate complex features through a cross-model modeling approach. Furthermore, CC-Time achieves the best prediction accuracy on two different types of datasets, demonstrating that this cross-model modeling approach can better adapt to diverse time series.
@@ 5 Conclusion
In this paper, we propose Cross-Model and Cross-Modality time series forecasting, namely CC-Time, to comprehensively model time series patterns. To leverage the capability of LLM for modeling complex patterns, CC-Time incorporates cross-modality modeling to capture temporal dependency and channel correlations in the LLMs from both time series sequences and their corresponding channel text descriptions. Furthermore, CC-Time proposes the cross-model fusion block to adaptively integrate knowledge from the LLMs and time-series-specific models to form a more comprehensive modeling of time series patterns. These innovative designs empower CC-Time to achieve state-of-the-art prediction accuracy in both full-data training and few-shot learning situations. At the same time, CC-Time exhibits better efficiency compared with most LLM-based methods.
@@ Appendix A Channel Descriptions
As shown in Figure 6, for generating semantic descriptions, we construct prompts for each channel: This is <dataset name>from <domain>, including <channel name1>, <channel name2>, …, please describe these channels and their correlations. We just need to input the dataset name, domain name, and the specific channel names into the predefined prompt. Then, by entering it into a large language model like ChatGPT, the semantic descriptions of channels are generated. For example, in the Weather dataset Wu et al. (2021), the description of the temperature channel (denoted as ”T (degC)” in the dataset) is generated from LLMs as follows: Temperature is a key parameter that describes climate conditions. It is closely related to humidity. As the temperature rises, the air can hold more moisture, affecting humidity. CC-Time can leverage this semantic information to understand the correlation between temperature and humidity in the real world, which helps in better modeling patterns.
@@ Appendix B Datasets
ETT (Electric Transformer Temperature) Zhou et al. (2021) collected from two different electric transformers, spans from July 2016 to July 2018, and includes 7 channels. ETT is divided into four subsets: ETTh1 and ETTh2, recorded hourly, and ETTm1 and ETTm2, recorded every 15 minutes. (2) Weather Wu et al. (2021) includes 21 different meteorological indicators that provide comprehensive data on weather conditions. These indicators, such as temperature, barometric pressure, humidity and others, offer a broad perspective on the atmospheric environment. (3) Electricity Trindade (2015) contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. (4) Traffic Wu et al. (2021) contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. (5) ZafNoo Poyatos et al. (2020) is from the Sapflux data project including sap flow measurements and environmental variables. (6) CzeLan Poyatos et al. (2020) is from the Sapflux data project including sap flow measurements and environmental variables. We split each evaluation dataset into train-validation-test sets and detailed descriptions of evaluation datasets are shown in Table 4. According to the channel description generation method proposed in section 3.2, the corresponding text for datasets is generated without requiring any additional human effort.
@@ Appendix C Baselines
  • S2IP-LLM Pan et al. (2024) aligns the semantic space of pre-trained PLMs with time series embedding space and enhances the representation of time series using semantic space informed prompting to improve forecasting performance.
  • FSCA Hu et al. (2025) builds the consistent context through structural alignment, logical alignment, and dual-scale GNNs, enabling PLMs to better understand time series.
  • Time-LLM Jin et al. (2024a) aligns the time series features into the language feature space through reprogramming techniques, then concatenates them with the text prompt before inputting them into a pre-trained large language model for feature extraction.
  • UniTime Liu et al. (2024b) designs domain instructions as prompts and uses the language model architecture for multi-source time series pre-training to extract broad knowledge.
  • GPT4TS Zhou et al. (2023) uses the pre-trained GPT2 as the backbone and adapts it to the time series space for feature extraction by fine-tuning the positional encoding and normalization layers.
  • iTransformer Liu et al. (2024c) treats the entire set of variables as tokens and uses the Transformer architecture to model the correlations between the entire set of channels.
  • Crossformer Zhang & Yan (2023) proposes a two-stage attention mechanism (cross-time attention and cross-dimension attention) to model the dynamics of time and the correlations between channels.
  • PatchTST Nie et al. (2023) employs a patching strategy to preserve local information and uses the Transformer architecture to capture the correlation between different patches for temporal modeling. It also applies a channel-independent strategy on the channel dimension.
  • FEDformer Zhou et al. (2022) proposes a frequency-enhanced decomposed Transformer architecture to model temporal dynamics from the perspective of frequency.
  • TimesNet Wu et al. (2023) transforms one-dimensional time series into a two-dimensional structure using Fourier transforms, enabling the modeling of 2D temporal variations to capture multi-periodicity time series patterns.
@@ Appendix D Compared with Time Series Foundation Models
To further validate the prediction performance of CC-Time, we conduct comparative experiments with several state-of-the-art time series foundation models, including Chronos Ansari et al. (2024), MORIAI Woo et al. (2024), Timer Liu et al. (2024e), and TTMs Ekambaram et al. (2024). The evaluation is performed on ETT and Weather datasets, with input lengths of $\{96,336,512\}$ and output lengths spanning $\{96,192,336,720\}$. The final results are selected based on the optimal performance across the three input lengths. As shown in Table 5, CC-Time achieves state-of-the-art prediction performance in most forecasting scenarios, outperforming the four baseline models. This demonstrates that CC-Time not only exhibits strong generalization capabilities but also effectively leverages knowledge from pre-trained language models, leading to significant improvements in time series forecasting accuracy.
@@ F.1 Effectiveness of Pre-trained Language Models
In the second experiment, refer to Tan et al. (2024), we further validate the effectiveness of PLMs in CC-Time. Specifically, we focus on three aspects: removing the entire PLM layers of CC-Time (W/O PLM), replacing the PLM layers with a single layer of untrained Attention (LLM2Attn), and replacing the PLM layers with a single layer of untrained Transformer (LLM2Trsf). As shown in Table 7, CC-Time effectively utilizes LLM knowledge to model time series. When comparing W/O PLM with CC-Time, the results indicate that relying solely on basic time series modules like Patching and ReVIN is insufficient, and that effectively exploring the potential of PLM knowledge for time series forecasting is crucial. When comparing LLM2Attn and LLM2Trf with CC-Time, the results indicate that relying on single-layer attention or transformer for feature extraction is inadequate, and that fully leveraging the strengths of PLMs and time-series-specific models to capture time series features comprehensively is essential.